### v1

# Edge AI: A Taxonomy, Systematic Review and Future Directions

Abstract
Edge Artificial Intelligence (AI) incorporates a network of interconnected systems and devices that receive, cache, process, and analyse data in close communication with the location where the data is captured with AI technology. Recent advancements in AI efficiency, the widespread use of Internet of Things (IoT) devices, and the emergence of edge computing have unlocked the enormous scope of Edge AI. The goal of Edge AI is to optimize data processing efficiency and velocity while ensuring data confidentiality and integrity. Despite being a relatively new field of research, spanning from 2014 to the present, it has shown significant and rapid development over the last five years. In this article, we present a systematic literature review for Edge AI to discuss the existing research, recent advancements, and future research directions. We created a collaborative edge AI learning system for cloud and edge computing analysis, including an in-depth study of the architectures that facilitate this mechanism. The taxonomy for Edge AI facilitates the classification and configuration of Edge AI systems while also examining its potential influence across many fields through compassing infrastructure, cloud computing, fog computing, services, use cases, ML and deep learning, and resource management. This study highlights the significance of Edge AI in processing real-time data at the edge of the network. Additionally, it emphasizes the research challenges encountered by Edge AI systems, including constraints on resources, vulnerabilities to security threats, and problems with scalability. Finally, this study highlights the potential future research directions that aim to address the current limitations of Edge AI by providing innovative solutions.
Index Terms:
Edge Computing, Artificial Intelligence, Cloud Computing, Machine Learning, Edge AI
I Introduction
Recent advancements in Artificial Intelligence’s (AI) effectiveness, the adoption of Internet of Things (IoT) devices, and edge computing capabilities are coming together to unleash the potential of edge artificial intelligence (Edge AI) [1]. Numerous analysts and businesses are conversing about and executing edge computing, which delineates its origins to the 1990s when edge servers positioned near customers were used to serve web and video content over content delivery networks [2]. In this edge AI, Edge computing is a paradigm transformation that brings data storage and processing closer to the data source, improving response times and reducing bandwidth usage. Unlike traditional cloud computing, where centralized data centres process data, edge computing processes data at the network’s edge [3]. This proximity reduces latency, enhances real-time data processing capabilities, and supports the expansion of IoT devices and services[4]. The primary advantages of edge computing include improved agility of services, low latency, enhanced coherence, and the elimination of a single point of failure, making it highly relevant for applications in smart cities, self-sufficient vehicles, and industrial automation [5]. By distributing computing resources geographically, edge computing ensures that data processing takes place near the data source, satisfying the need for analytics and decision-making in real time.
On the other hand, AI includes a wide array of technologies and methodologies that enable machines to carry out tasks that generally require human intelligence, such as learning, reasoning, and self-correction [6]. AI’s applications span various domains, including healthcare, finance, transportation, etc, where it is used to analyze large datasets, automate tasks, and provide predictive insights [7]. Integrating AI into different sectors has revolutionized processes by enhancing efficiency, improving decision-making, and creating new opportunities for innovation. With betterment in Machine Learning (ML) or Deep Learning (DL), AI approaches have become increasingly competent in performing complex tasks that require human-like cognitive functions [8]. AI algorithms, specifically those involving neural networks, have shown remarkable success in areas like image and speech recognition, autonomous driving, and predictive maintenance.
I-A Edge Computing for AI
The fusion of edge computing and AI involves processing AI algorithms on users’ devices, offering benefits like reduced latency, energy efficiency, and real-time applications. This integration allows for real-time data processing and decision-making at the source, significantly declining latency and bandwidth use [9]. The combination of edge computing and AI enables the development of smarter and more responsive applications, such as autonomous vehicles, industrial IoT, smart home systems, etc. By leveraging edge AI, organizations can achieve greater efficiency, enhanced privacy, and faster insights, driving innovation across various sectors [10]. Edge AI refers to integrating AI capabilities at the network edge, enabling distributed intelligence with edge devices. It intends to improve network connectivity, enable deployment of AI pipelines with defined quality targets, and allow adaption for data-driven applications.[11]. Embedding AI functionalities at the edge addresses the limitations of cloud-based processing for IoT, such as privacy concerns and network connectivity issues. The deployment of AI at the edge enhances latency-sensitive tasks and reduces network congestion, improving efficiency and security in wireless networks.
Furthermore, AI-based technologies play a vital role in addressing Quality of Service (QoS)-aware scheduling and resource allocation challenges in edge environments, ensuring quality of service and user experience. Edge AI enables the deployment of AI as a Service (AIaaS) with configurable model complexity and data quality, enhancing performance and reducing costs [12, 13]. This innovative approach supports smart security applications by leveraging AI capabilities at the edge and enhancing security measures for distributed systems. Edge intelligence, a promising technology, empowers real-time applications by moving computing from cloud servers to IoT edge devices, creating intelligent enterprises with vast possibilities [14]. The utilization of AI at the edge, instead of centralized locations, unlocks the potential of AI with IoT devices and edge computing, deploying AI algorithms on resource-constrained edge devices for various applications like autonomous vehicles, healthcare, and surveillance.
Edge AI’s significance is underscored by its ability to provide immediate insights and actions without sending significant amounts of data to several centralized locations [15]. This capability is particularly critical in scenarios where latency and bandwidth are significant constraints, such as in autonomous driving, where decisions must be made in real time, or in healthcare, where patient data must be processed quickly to provide timely interventions [16]. The rise of edge AI is also fueled by advancements in hardware, such as more powerful and energy-efficient processors, which make it feasible to run sophisticated AI models on devices like smartphones and IoT sensors [17].
I-B Motivation and Contributions
The motivation for integrating edge computing with AI is multifaceted, primarily driven by the imperative need for processing data in real time and navigating the inherent limitations of centralized cloud computing systems [18]. As we witness an exponential rise in the number of connected devices and a corresponding surge in data volume, traditional cloud-centric models increasingly grapple with issues such as latency, bandwidth constraints, and significant data privacy concerns. Edge AI emerges as a pivotal solution to these challenges, advocating for localized data processing [19]. This shift not only diminishes the reliance on distant cloud infrastructures, thereby slashing latency, but also significantly bolsters the responsiveness of applications to real-time data inputs. This paradigm shift is particularly pivotal for fueling the development of next-gen technologies that necessitate instantaneous data analysis and decision-making, encompassing sectors like autonomous vehicles, smart city infrastructures, and cutting-edge healthcare systems.
Moreover, Edge AI empowers applications to operate remarkably efficiently, even in scenarios characterized by sparse connectivity, by facilitating data processing directly at the source. This capability is indispensable in remote or highly mobile environments where consistent and reliable internet access is only sometimes assured [20]. By processing data onsite, edge AI considerably amplifies data privacy and security measures, mitigating the need to transmit sensitive information over vast distances to central servers. This feature is exceptionally critical in domains such as healthcare and finance, where the confidentiality and integrity of data are of utmost concern.
Additionally, Edge AI champions bandwidth efficiency by mitigating the volume of data that needs to be transmitted over networks, making it an economical choice for data-intensive applications [21]. This efficiency not only reduces operational costs but also relieves network congestion, facilitating smoother and more reliable data flows. Scalability is another significant advantage offered by edge AI [6]. As the network of devices expands, edge computing allows for seamless scalability without the bottleneck of centralized processing power, enabling businesses and technologies to grow without being hampered by infrastructure limitations.
Essentially, the combined use of edge computing and AI is not just a technical progression but also a tactical imperative to fulfill the dynamic requirements of contemporary applications. By championing lowered latency, enhanced privacy and security, bandwidth efficiency, and scalability, edge AI is set to revolutionize how data-driven decisions are made, ushering in a new era of intelligence that is both efficient and privacy-centric.
The main contributions of this paper are:
  • We offer a thorough introduction to Edge AI, covering its history, challenges, and prospects.
  • We conducted a systematic review that provided a thorough examination of edge AI research based on many applications, highlighting current trends and possible directions for the future.
  • We propose a taxonomy for edge AI, which aids in the classification and arrangement of edge AI systems, and explore its potential impact across disciplines through various applications.
  • We emphasize how important edge AI is for processing data in real time at the network’s edge. It also highlights the challenges faced by edge AI systems, such as resource limitations, security risks, and scaling issues.
  • We propose promising future directions that aim to address the current shortcomings of Edge AI by providing innovative solutions and opportunities for future research.
I-C Article Organization
Section II discusses a related survey focusing on different applications in terms of algorithms, optimization techniques, security, and privacy concerns integrated with Edge AI. Section III covers the background and present present situation of Edge AI. Section IV details the methodology adopted for the review. Section V outlines a taxonomy encompassing infrastructure, cloud computing, fog computing, services, use cases, machine learning and deep learning, and resource management. Section VI compares existing Edge AI implementations based on taxonomy. Section VII presents an analysis and the results obtained, and the future research directions are discussed in Section VIII. Finally, Section IX summarizes the survey.
II Related Surveys
In this section, we discuss various studies, which are about the different applications consisting of smart cities, smart manufacturing, autonomous vehicles, the Internet of Vehicles (IoV), industrial automation, and healthcare monitoring systems. These are highlighted when edge computing meets AI and AI for edge computing. There are also considerations of traditional ML, computation offloading optimization, and concerns related to privacy and security, reflecting a comprehensive analysis of the challenges and strategies integrating AI and edge computing.
II-A Smart Cities
In the case of innovative city applications, The intersection of AI and edge computing in smart cities emphasizes the importance of optimizing computation offloading and fostering a Federation between edge, cloud, and fog computing for efficient operations, which have been discussed in the following articles. In 2020, authors [22] presented an intelligent offloading method (IOM) that preserves privacy, boosts edge utility, and improves offloading efficiency for smart cities. The mechanism of information entropy is utilized in conjunction with edge computing to achieve an equilibrium between the maintenance of privacy and the facilitation of collaborative services. Further, authors [23] employed a cooperative compute offloading method to obtain the aforementioned trade-off in the cooperation of three ends: IoT device, cloudlet, and cloud. Offloading, on the other hand, can significantly reduce the processing strain on IoT devices; yet, it may incur high transmission costs and cloudlet resource use. Furthermore, authors [24] outlined the introduction of a cyclic branch network, a DL-based intelligent offloading scheme that makes full use of network edge computing power and data traffic to reduce overall energy consumption in dual connectivity and nonorthogonal multiple access computation offloading systems. Moreover, authors [25] suggested Robust Neural Networks From Coded Classification (CoDNN), a unique compute offloading method for multi-device collaborative pipelining processing of deep neural network (DNN) tasks. In [26], authors focused on an approximation technique called Accuracy Maximization using LP-Relaxation and Rounding (AMR2), which is suggested and shown to produce a makespan of no more than 2T and a total accuracy that is less than the optimal total accuracy by a small constant. Another work [27] presents a revolutionary deep neural network-based energy-efficient offloading strategy that trains a smart decision-making model to select a reliable pool of application components. Finally, researchers [28] suggested method for the heterogeneous scenario is consistently effective in identifying a superior offloading scheme than the chosen existing algorithms, according to empirical findings. On the other hand, for the homogeneous scenario, the suggested solution can effectively accomplish the ideal approach.
II-B Smart Manufacturing
In intelligent manufacturing, combining AI with Edge Computing significantly enhances efficiency, decision-making, and security, effectively addresses challenges, and improves data utilization. This integration streamlines operations, enables predictive maintenance, and supports autonomous decisions. The following papers are discussed to describe these advancements and subsequent challenges: In 2022, authors [29] explored the scope of AI and its application in India’s intelligent manufacturing industry, concentrating on the technology’s current state, constraints, and recommendations for resolving issues. In [30], authors have talked about how ML and AI may boost productivity, sustainability, and manufacturing efficiency. However, there are a number of difficulties with implementing AI in manufacturing, including problems with infrastructure and human resources, security threats, trust, and data management and acquisition. Further, the authors [31] suggested a new mode called ”AI-Mfg-Ops” (AI-enabled Manufacturing Operations) with a supporting software-defined framework is proposed as part of an open evolutionary architecture of the intelligent cloud manufacturing system. This mode can facilitate quick operation and upgrades of cloud manufacturing systems with intelligent assessment, analysis, planning, and execution in a closed loop. In 2021, authors [32] addressed the job shop scheduling challenge in the intelligent factory process while using a Deep Q-network (DQN). The suggested framework is contrasted and examined with other frameworks from the standpoint of offering an intelligent factory service. Furthermore, the authors [33] look at how it tends to integrate several productivity factors, such as big data analytics, Automation, and Operations Information, which connect machines via open platforms, resulting in real-time reactions to boost efficiency across the supply chain. Moreover, the authors[34] developed a service-oriented information model to standardized describe the functional characteristics and related operational data of heterogeneous manufacturing resources; additionally, a message middleware-based real-time transmission and integration method for high-volume operational field and sensor data is suggested in order to achieve the efficient distribution of related data and remote monitoring of distributed manufacturing resources.Fianlly, the authors [35] discusses the possible advantages and difficulties of a federated learning architecture based on data gathered from 5G MSPs to enable predictive maintenance (PM) in industrial settings.
II-C Autonomous Vehicles and the Internet of Vehicles (IoV)
In the context of autonomous and IoT-enabled vehicles, advancements in control and task optimization are propelled by AI integration and Edge Computing (EC). This synergy supports real-time decision-making, highlighting the importance of AI and EC in addressing challenges like real-time processing and security and optimizing communication and privacy within the IoV. The application of DL and Reinforcement Learning with Multiple Agents (MARL) Underscores the need for efficient solutions in autonomous driving technologies. In these scenarios, the following papers are described: In 2023, authors [36] offers a thorough technical overview of the most recent studies conducted in the areas of lateral, longitudinal, and integrated control strategies for self-driving cars. They also examine a variety of strategies and tactics used to attain accurate steering control while taking longitudinal factors into account. [37] discusses key technologies, applications, solutions, and problems related to integrating Mobile Edge Computing (MEC) and ML in the Internet of UAVs and are covered in-depth by the author’s thorough review. Further, authors [38] examined the most recent research on vehicular data offloading from the standpoint of communication, focusing on vehicle-to-vehicle (V2V), vehicle-to-roadside infrastructure (V2I), and vehicle-to-everything (V2X). The study also identified unresolved research issues in this area and forecasted future directions in the field. Furthermore, the authors [39] suggested a multi-access edge computing (MEC) framework to facilitate the cooperation of digital twins (DTs) into wireless networks and connected cars (CVs) to reduce the unreliability of long-distance communication between edge servers and CVs. Moreover, researchers [40] outlines a DL and edge computing-based vehicle intelligent control system that encourages the broad advancement of automation and intelligent technology. The findings show that the distance between the target and experimental vehicles is extremely close to the anticipated safe space. In [41], authors suggested a safe edge intelligence that combined the advantages of blockchain, local differential privacy (LDP), and federated learning (FL) for automotive networks. The authors in [42] focuses on a fast task execution technique in heterogeneous IoT applications that is powered by AI. This technique reduces decision latency by considering various system parameters, including the task’s execution deadline, the device’s battery level, the channel conditions between mobile devices and edge servers, and the capacity of the edge servers. Finally, the authors [43] combined edge computing and the Web of Things, compared their functions, and showed how edge computing improves the efficiency of real-time IoT applications by focusing on transmission, storage, and computation elements.
II-D Industrial Automation
In industrial automation, several papers discuss revolutionary approaches to enhancing productivity by integrating AI, edge computing, robotics, and data analytics. The relevant papers look over the utilization of the Industrial Cyber Intelligent Control Operating System (ICICICOS), a cloud-edge computing-based system, for AI and industrial automation. It focuses on proposing AI with industrial processes at the edge. It emphasizes strategies for optimizing ML methods, deploying AI models on resource-constrained devices, and addressing security concerns through secure AI microservices at the edge. The relevant papers are described as follows: In 2022, the authors [11] presented a flexible working mechanism by permitting the combined design of data quality ratios (DQRs) and model complexity ratios (MCRs) for the AI tasks and suggested a configurable model deployment architecture for edge AIaaS. Further, the authors [12] provide a systematic overview of methods for addressing the paucity of training data for different kinds of data, and a methodology for addressing data scarcity in cellular networks is suggested. In [14], the authors explore the privacy-enhancing solutions that are now in position, including the technologies, specifications, and process solutions to mitigate these risks. It also looks at privacy threats at various stages of the AI life cycle. Further, the authors [16] offer a thorough overview of edge intelligence and lightweight ML support for upcoming services and applications. The researchers have supplied a thorough analysis of cutting-edge intelligence applications, lightweight ML techniques, and their support for upcoming services and applications. In [44], authors provide a thorough review of AI/ML-based IDS/MDSs and set baseline measurements pertinent to networked autonomous systems, emphasizing the gaps and assessment metrics in the existing research. In [6], the authors wrapped up a thorough analysis of edge computing, covering both the shift to edge AI and related paradigms. Additionally, the history of every alternative put out for edge computing implementation, as well as the Edge AI strategy for putting AI models and algorithms on edge devices, were investigated.
II-E Smart Healthcare
Intelligent healthcare systems focus on integrating AI and edge computing and the challenges related to privacy and security, decision-making, and optimization. These systems make use of technologies, including genetic-based encryption for data security, federated learning in the Internet of Medical Things, and nanosensor-equipped systems to improve efficiency and security. Mobile computing has played a vital role in healthcare, particularly during the COVID-19 pandemic, enabling telemedicine and contact tracing, emphasizing the significance of technology in tackling healthcare challenges. The following research papers delve deeper into the intersection of technology, healthcare, and privacy. In 2018, the authors [45] proposed an intelligent home monitoring system based on edge-fog computing with AI capabilities. Latency issues and reliability are the main concerns for the authors in developing smart home real-time applications. Further, the authors [46] provide a comprehensive overview of the key elements of the MCPS from multiple perspectives, covering design, methodology, and significant supporting technologies such as cloud computing, edge computing, IoT, sensor networks, and systems with multiple agents. In [47], authors provided a distinctive and specialized route resource recommendation (R3) protocol to handle resource management and connection problems in autonomous, connected ambulances (ACA) for route optimization. Furthermore, the authors [48] provide a condition-aware analytical framework that may be used to recommend health conditions in the IoT-based mobile healthcare systems. This framework corresponds to IoT devices that have limited resources, such as those with a memory utilization rate of 6.6%. Moreover, the authors [11] Present a flexible working mechanism that allows the combined configuration of data quality ratios (DQRs) and model complexity ratios (MCRs) for AI tasks and addresses a flexible model deployment architecture for edge AIaaS. In [49], researchers suggest using an edge-of-things (EoT) framework to implement centralized and federated transfer learning (CMTL) for cyberattack detection systems in the healthcare industry. Finally, the authors [50] presents a new approach called CoDoC, which stands for complementary-driven deferral-to-clinical workflow. Its purpose is to decide when to rely on a diagnostic AI model and when to hand it off to a clinician.
TABLE I: Comparison of Our Survey with Existing Related Surveys
Work Research Domain [2] [4] [6] [7] [8] [9] [10] [11] [12] [14] [16] Proposed work
Year 2022 2023 2023 2020 2022 2024 2020 2023 2024 2023 2023 2024
Edge AI ✓(*) ✓(*) ✓(*) ✓(*)
Taxonomy ✓(*) ✓(*)
Infrastructure Cloud
Fog
Edge ✓(*)
Application Monolithic ✓(*)
Architecture Microservice ✓(*)
IoT Static ✓(*)
Use Cases Mobile ✓(*)
Methods Heuristic
Meta-Heuristic
Machine ✓(*) ✓(*)
Learning
Deep Reinforcement ✓(*)
Learning
Provisioning ✓(*)
Resource Resource Allocation ✓(*)
Management Application Placement ✓(*) ✓(*)
Workload Distribution
and Prediction
ML Model Sizing Reduced
Full
Heterogeneity Computational ✓(*) ✓(*) ✓(*)
Hardware
Platform
Security Platform
Host
Network ✓(*)
Scheduling Container ✓(*) ✓(*)
Task
Pod
Service
Container Migration Stateful vs Stateless containers
Inter versus Intra cluster migrations
Migrations at cloud/Edge/fog ✓(*) ✓(*) ✓(*)
Simulations versus real-world testbed migrations ✓(*)
Container Scaling Proactive versus reactive scaling decisions
Horizontal, Vertical and Hybrid scaling ✓(*)
Abbreviations: ✓:= method supports the property. *:= just an Overview/Visionary
II-F Comparison with Related Surveys
Table I shows the comparison of our systematic review with related surveys by focusing on the advancement of AI at the edge, optimizing algorithms in constrained environments, solving training data scarcity with AI techniques, and using AI/ML for resource management in fog and edge computing settings. There, the different columns are set on the basis of various applications, their methods of optimization, and fruitful utilization. This survey covers several edge computing concepts and emphasises the application of AI on edge devices with constrained resources. The survey examines the optimisation of ML algorithms for such restricted environments and covers current IoT applications across several sectors, including industrial automation, smart homes, and autonomous vehicles. It also highlights the difficulties and possible paths for edge computing and edge AI research, offering a strong basis for further investigation in the field. Considering the dynamic, diversified, and resource-constrained character of fog and edge computing settings, this research emphasises the potential of AI/ML, particularly reinforcement learning techniques, in addressing resource management challenges in these systems [2] [4] [6] [7]. A thorough analysis of the literature was done to look at how AI/ML applications may be used to effectively manage resources in these kinds of situations. A taxonomy was established to categorise and contrast different approaches. Enhancing explainability, reducing variance, and boosting online training of AI/ML algorithms are highlighted as critical future research directions to adapt to the constantly changing fog/edge computing landscapes. The study emphasises the significance and changing challenges of resource management [8], whereas a presentation framework is described that addresses the issue of sparse training data in emerging radio access networks by utilizing a range of methods, such as interpolation, domain-knowledge techniques, generative adversarial networks (GANs), transfer learning, autoencoders, few-shot learning, simulators, and testbeds. The challenges are highlighted and presented by insufficient training data, and the crucial role that Automation powered by AI plays in the operation, optimisation, and troubleshooting of cellular networks is described. The technique suggests an integrated strategy to improve data availability in cellular networks and includes a survey and taxonomy of current approaches to lessen this scarcity. In addition, the study emphasises the necessity for scalable, reliable solutions that take conditional contexts into account for generating high-dimensional data in radio access network applications also [9] [10]. Table I reveals a dearth of reviews on Edge AI, with most existing surveys and review papers presenting an overview or vision of the technology rather than a comprehensive survey, systematic review, and detailed taxonomy. To the best of the author’s knowledge, this is the first survey paper on edge AI that provides a thorough taxonomy and a systematic review and highlights future research topics.
III Edge AI: Background and Current Status
This section explains some concepts related to background and current status in Edge AI for the reader. Subsection III-A explains edge computing and its historical emergence. Subsection III-B provides information on the integration of AI and edge technologies. This section is completed by explaining Edge AI applications and challenges in subsection III-C and subsection III-D, respectively.
III-A Historical Emergence of Edge Computing
The concept of edge computing is a paradigm that brings computing resources closer to the data source, unlike the cloud, which provides services through a remote server [51]. In this way, it is aimed to reduce problems such as unnecessary bandwidth occupation and latency in today’s world where huge amounts of data that need to be processed are produced [52]. To understand the emergence of edge computing, it will be more useful to examine previous paradigms such as cloud and fog computing. Figure 1 shows the advantages of cloud, fog, and edge computing over each other and their layer arrangement. Now let’s briefly touch on these concepts:
Figure 1: Computing Paradigms and their Objectives.
  • Cloud Computing: Cloud computing is a paradigm that dates back to the 1970s and refers to the use of common computing resources by users on a server via the Internet [53]. Today, it is offered to users with various service models, especially by large companies such as Microsoft Azure, Google Cloud Platform and IBM Cloud. The advantages of cloud computing are as follows [54]:
    • High processing power and central storage, so users can easily access resources from anywhere there is the internet. This reduces the user’s risk of data loss and provides the user with the freedom to work from any location with internet access.
    • Scalability, in case the need for computing resources increases (demand fluctuations), cloud computing provides services such as more processing power and storage by scaling the resources. In this way, performance measures such as SLA and QoS are ensured.
    • Pay as you go, with the serverless (Function as a Service (FaaS) + Backend as a service (BaaS)) service model provided by cloud computing, users are charged only for the amount they use their computing resources. In this way, an economical model is provided and appealed to more users.
  • Fog Computing: The concept of fog computing was introduced by Cisco in 2012 [55]. This paradigm recommends moving computing resources closer to the endpoints of the network (such as routers and gateways) to reduce the latency and bandwidth problems that occur in cloud computing. When Figure 1 is examined, fog computing acts as a layer between the cloud and the edge. The advantages of fog computing are as follows [56]:
    • Fog computing has lower latency than cloud because it brings computing resources closer to the edge of the network.
    • By acting as a layer between the cloud and end devices, it reduces unnecessary bandwidth usage by processing some of the huge amounts of data to be sent to the cloud.
  • Edge Computing The development of IoT and sensor technologies has increased the amount of data that needs to be processed to enormous levels. Processing all this data on cloud computing resources may cause unnecessary bandwidth occupation and latency problems. For this reason, the concept of edge computing has emerged as a paradigm that aims to optimize latency and bandwidth usage by processing data close to the data source [57]. Additionally, edge computing is a good solution to address the complexity, security, and management challenges posed by fog computing, an extra layer [58]. The advantages of edge computing are as follows [59]:
    • Reduces latency and bandwidth usage by moving data processing to the edge of the network,
    • Compared to fog computing, it offers advantages such as less complexity and better security.
III-B Integration of AI with Edge Technology
The idea of combining the concepts of AI and edge computing is one of the hottest topics of recent times. Fig. 2 shows a main diagram explaining the concept of Edge AI. Edge AI is processing data in AI algorithms distributed on edge nodes (IoT and mobile devices) close to the data source, rather than processing data on a central cloud platform [60]. The advantages offered by these two technologies can be listed as follows [6]:
  • The amount of latency is critical in applications such as low latency, time-sensitive autonomous vehicles, and e-health where patients are monitored. With Edge and AI integration, this problem can be overcome by processing data in real time.
  • In the data security, edge paradigm, since the data will be processed at a point close to the source, sensitive data such as biometric data is processed on local devices within the network. In this way, security possibilities such as data leakage can be eliminated.
  • Scalable resources, edge paradigm consists of nodes consisting of homogeneous or heterogeneous devices. In this way, task scheduling can be made according to processing power needs.
Figure 2: Architecture of Edge AI.
III-C Edge AI Applications
Edge AI applications, created by combining the concepts of Edge and AI, provide lower latency and higher security than Cloud-based AI applications. Figure 3 shows some of these applications. Let’s explain these applications:
  • Healthcare: Edge AI applications are based on the processing of data collected from wearable devices in distributed AI models at the edge of the network. Additionally, early diagnosis studies using portable medical imaging techniques can be given as examples [54].
  • Smart Parking: With the increase in means of transportation, parking has become a big problem, especially in big cities. Edge AI-based solutions with the help of sensors and IoTs can be used to solve these problems [61].
  • Smart Home: Solutions used in modern homes such as home lighting systems and smart refrigerators can be given as examples of these applications. In this way, energy consumption can be optimized by preventing excess electricity consumption in cities [56].
  • Computer Vision: Edge AI can identify people using methods such as biometric authentication [51]. Additionally, Edge AI provides great advantages in Industry applications that require real-time decisions [58].
  • Cyber Security: Unauthorized access, suspicious objects, and armed individuals can be detected with Edge AI-based security applications. Additionally, anomaly detection can be made by detecting suspicious traffic on a network to prevent cyber attacks [62].
  • Transportation: Edge AI-based solutions can be used for today’s complex traffic light operations [63].
Figure 3: The Edge AI Applications.
Figure 4: Edge AI Implementation Challenges.
III-D Edge AI Implementation Challenges
EdgeAI, which emerges by combining Edge and AI, brings with it the advantages it offers, but also challenges that are still waiting to be solved. These challenges are shown in Figure 4. Let’s explain these challenges:
  • Energy Efficiency: Edge devices generally consist of homogeneous and heterogeneous devices with low processing and storage capacity. Applications that require Natural Language Processing (NLP) and intensive image processing will cause excessive resource consumption on edge devices [64]. For this reason, new solutions such as special AI chips or task engineering are needed.
  • Maintenance and Updates: Since edge devices consist of devices distributed in different locations, this means more attack targets for hackers [65]. In addition, not all devices in the edge nodes have a homogeneous structure, which means separate system maintenance and updates for each node [66]. Measures such as automatic updating can be taken to solve these problems.
  • Scalability: Since edge devices generally consist of heterogeneous devices, the distribution of a single application to different devices is still a challenge (task scheduling, etc.) [67]. Additionally, it is difficult to synchronize data across all devices. Effective microservice architectures and load-balancing algorithms that prevent a node from being overloaded can be used to solve this problem.
IV Review Methodology
This article does a systematic review to categorise studies that are pertinent to this study domain or discuss specific research questions on ”Edge AI”. In this article, we used the guidelines established by Kitchenham et al. [68, 69, 70] to provide a comprehensive review of Edge AI. The review is the optimal and reliable approach to document and analyse current research works. The systematic approach enables researchers to carefully analyse the positive and negative aspects of recent studies, conduct a thorough examination to identify potential gaps in research and future trends and difficulties, and provide a solid foundation and context for establishing a new study field. Furthermore, the complete research approach is presented in Figure 5, that represents the structure of the process that was used in the systematic review research.
Figure 5: Road-map for Research Methodology
IV-A Design and Plan of Review
The review procedure illustrates the methodologies applied to conduct a systematic review with the aim of minimizing the potential for biased research. Therefore, possessing a pre-established process is crucial. In the absence of a systematic methodology, researchers’ predispositions have the ability to influence the process of selecting and analyzing studies. This may result in the omission of crucial inquiries essential for a thorough analysis and comprehension of the subject matter. The review process encompasses the research inquiries, exploration approach, criteria for selecting studies, procedures for assessing quality, and techniques for extracting and synthesizing data [71].
IV-B Research Questions
Determining the research queries is crucial in the method of planning to develop a strong systematic review. The design of the research problem requires a thorough examination of existing literature studies. The primary aim of the present systematic review is to thoroughly examine and evaluate the various methods and strategies being employed for the  edge intelligence or edge AI. Furthermore, in order to emphasise the research findings and effectively showcase the useful consequences, the research questions that follow have been defined in Table II.
TABLE II: Research Questions, Motivation, Category, and Mapping
Sr. No. Research Question Motivation Category Mapping Section
RQ1 What are the primary methodologies and strategies adopted in Edge AI, and how do they vary in terms of applications, infrastructure, and algorithms? The purpose of this research question is to understand the various methods and approaches used in Edge AI. Infrastructure, Application Architecture, and Methods Sections 5.1, 5.2, and 5.4
RQ2 To what extent may heuristic algorithms enhance the effectiveness and performance of Edge AI systems, specifically in tasks such as model training and optimization? This research question helps in examining the role of heuristic methods in Edge AI. Heuristic Methods Section 5.4.1
RQ3 How do the meta-heuristic algorithms used by Edge AI solve difficult optimization problems, and what are their applications and prime characteristics? The characteristics and applications of meta-heuristics methods are investigated in this paper. Meta-Heuristics Methods Section 5.4.2
RQ4 How can real-time data processing and decision-making in Edge AI devices get aided by ML techniques, including deep learning models? This research question analyzes the role of ML techniques. Machine Learning Methods Section 5.4.3
RQ5 What are the main benefits and constraints of Edge AI’s deep reinforcement learning, specifically for situations like autonomous driving? The pros and cons of deep reinforcement learning are examined via this research question. Deep Reinforcement Learning Methods Section 5.4.4
RQ6 How do Edge AI devices deal with heterogeneity in computational, hardware, and platform factors, and what approaches are used to optimize resource allocation? Heterogeneity in various aspects is discussed along with optimization of resource allocation approaches. Heterogeneity and Horizontal, Vertical, and Hybrid container scaling Sections 5.7 and 5.11.2
RQ7 What are the main challenges and measures for implementing Edge AI applications, specifically in terms of resource provisioning and workload distribution? The purpose of this research question was to identify challenges and solutions while implementing Edge AI applications. Resource Provisioning and Workload Distribution and Prediction Sections 5.5.1 and 5.5.4
RQ8 What are the pros and cons of both monolithic and microservices architectures in terms of Edge AI flexibility, performance, and resource utilization? This research question helps in comparing both monolithic and microservice architectures. Comparisons of Existing Edge AI approaches on the basis of Application Section 6.2
RQ9 What are the primary factors worth considering for resource management in Edge AI including resource provisioning, allocation, application placement, workload distribution, and prediction? The prime factors for resource management are examined through this research question. Resource Management Section 5.5
RQ10 How do ML model sizing impact the performance and resource consumption of Edge AI applications? This research question investigates the impact of machine learning model size on the performance and resource consumption of Edge AI applications. Full ML Model Sizing Section 5.6.2
RQ11 What are the key components that affect the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, with regard to how they affect application performance and resource utilization? This research question helps identify components that influence the choice of Edge AI infrastructure. Infrastructure Section 5.1
IV-C Search Strategy
IV-C1 Database Selection
The database selection includes conducting searches on various digital databases, such as IEEE Xplore, ACM Digital Library, Wiley, Taylor and Francis, Springer Link, Google Scholar, and Science Direct. These databases contain a wide range of impact factor journals, magazines, and significant conference proceedings, making them suitable for this systematic review.
IV-C2 Search Query
A comprehensive search was conducted utilising Logical OR/AND operators to connect the keywords, concepts, synonyms, and abbreviations. The initial phase entails conducting an automated search using predetermined keywords that align with the study topics of this systematic literature review (SLR). The keywords used are [(((”Edge AI” OR ”Edge Intelligence” OR ”Edge Computing”) AND (”Machine Learning” OR ”Deep Learning” OR ”Reinforcement Learning”)) AND (”Resource Management” OR “Resource Allocation”)) OR (“Cloud Computing” OR “Fog Computing”) OR (“Application Placement” OR “Security” OR “Scheduling” OR “Simulation”)]. The search terms are obtained from the specified research topics and the framework of this systematic review in order to encompass the most significant and interrelated publications.
TABLE III: Overview of the criteria for determining inclusion and exclusion.
S. No. Inclusion Exclusion
1 English articles issued at conferences, journals, and book chapters. Non-English articles.
2 Articles that are included in a database source and are available in their entirety. Articles that are not available in their whole
3 Articles that specifically examine the process of choosing Edge AI infrastructure, such as Cloud, Fog, and Edge computing, and their effects on application efficiency and consumed resources. Articles that explore diverse domains such as federated learning, IoT-based approaches, and other classic methods.
4 Articles published till 2024. Articles that were not published during the designated search timeframe.
5 Relevant articles pertaining to the investigation queries. Articles that fail to meet the research requirements or receive a score of 3.5 or lower in the quality assessment standards.
6 Systematic reviews often prioritise publications containing experimental or empirical research. Articles that do not contain such research.
IV-D Inclusion and Exclusion Criterion
The systematic review needs to establish clear guidelines for inclusion and exclusion to guarantee that the chosen papers are relevant to the research topic and address the specific research objectives. The primary purpose of establishing the criteria is to guarantee that the studies featured are appropriate and connected to AI-based methodologies in Edge computing. Hence, the chosen research must satisfy all the predetermined criteria. Table III presents the specific phrases used to determine which criteria were included and excluded in this systematic review.
Furthermore, a method of screening is carried out to identify the appropriate research studies that are relevant to the context of this study shown in Fig. 6. The screening process consists of three distinct stages:
  1. a.
    Title and abstract Phase: During this stage, papers that were deemed irrelevant were excluded based on their title and abstract. Subsequently, the studies that satisfy at least some of the criteria listed in Table III are chosen and advanced to the subsequent step for additional analysis.
  2. b.
    Full-text screening Phase: During this step, studies were excluded while they failed to fulfill the criteria specified in Table III, based on a thorough reading of the full-text or partial reading.
  3. c.
    Final selection Phase: The next phase utilises the criteria terms outlined in Fig. 6 to make the final selection and eliminates studies that do not meet any of the specified criteria.
    Figure 6: Research Methodology Protocol
    1. i.
      The topic of the study must be pertinent and directly connected to the research topics.
    2. ii.
      The user did not provide any text. The research study examines the comprehensive solution for research advancements in edge intelligence and identifies four key components: monolithic and microservices architectures differ in terms of flexibility, performance, and resource utilisation in Edge AI for addressing practical challenges, finding solutions, and achieving optimisation goals.
    3. iii.
      The research paper presents essential factors to consider for managing resources in Edge AI, such as resource provisioning, allocation, deployment, and scheduling of workloads.
    4. iv.
      The user did not provide any text. The research study examines the factors that influence the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, and investigates their effects on application reliability and resource utilisation.
    5. v.
      The user did not provide any text. How do Edge AI systems maintain reliable and intelligent tasks in dynamic and ambiguous instances?
IV-E Quality Assessment
In order to gather the most comprehensive and reliable information on this subject, we employed a systematic review methodology, following the standards outlined [72]. Furthermore, a plethora of research papers and conference papers exist on the topic of AI applied to edge computing. Once we applied the criteria for inclusion/exclusion, researchers conducted a thorough evaluation of the articles that met the standards to identify the ones that were most worthy of further examination. Employed the standards established by the methodology to evaluate the overall quality of the research, taking into account its impartiality, internal consistency, and objectivity.
IV-F Extraction and Synthesis
This phase emphasises the process of extracting and combining data by thoroughly examining all 78 chosen studies and summarising and storing the relevant information. This stage involves the creation of a mechanism for extracting data items and compiling comprehensive reports that include all the information gathered from primary research studies [71].
Furthermore, this study specifically chose items that relate to the research objectives as well as research questions. The data was taken from primary studies and meticulously recorded to determine the ultimate findings of the systematic review. The analysis step involved employing descriptive synthesis. The subsequent section discusses the results obtained from the synthesis.
V A Taxonomy of Edge AI
Figure 7: The Taxonomy of EdgeAI.
Based on the current advancements and existing literature on edge AI, we have proposed a taxonomy, as shown in Fig. 7.
V-A Infrastructure
AI models and applications can be deployed in different infrastructures based on their target scenarios. The three most common infrastructures widely used are Cloud, Fog and Edge.
V-A1 Cloud
The essence of cloud computing is to pool resources with virtualization technology. Virtualization technology transforms a physical machine into several virtual machines, greatly changing the mode of application operation and deployment. The term ’cloud’ here refers to remote data centers [21]. Users usually connect to the servers in the data center through the network to use the computing resources here. The essence of cloud computing is to pool resources, allowing users to purchase resources according to their own requirements, greatly reducing resource waste.
Therefore, through cloud computing, the performance of AI applications has also been significantly improved. We transmit local AI models to remote servers through the network. We can more efficiently utilize some computing resources in the cloud, such as CPUs, GPUs, and memory, and flexibly scale according to our business needs, improving resource utilization and reducing the cost of model training and prediction.
V-A2 Fog
As is well known, fog is closer to the ground than cloud, and the ground here refers to the user’s local device. Fog computing technology adopts distributed computing technology, which arranges computing resources on the side closer to the user than cloud computing [6]. It can be said that fog computing has broadened the network computing mode of cloud computing, widening computing capacity from the network center to the network edge, and thus more widely applied to numerous services. Fog computing is more widely distributed geographically and has greater mobility, making it suitable for an increasing number of intelligent devices that do not require extensive computation. For some time delay-sensitive applications such as real-time interaction system, fog computing also has greater advantages.
Image, video and natural language processing (NLP) are the recently emerging applications of fog computing. The placement and processing of images in fog computing is one of the most widely used fields of AI in research and industry. Its goal is to distinguish objects and people from each other and classify and distinguish photos based on image processing algorithms. Using fog computing in image processing-based applications can shorten response delay and improve service quality. In medical applications that require image processing accuracy and fast processing of medical data, deploying effective scheduling algorithms in foggy environments may be beneficial. According to other works, DL algorithms such as Generative Adversarial Networks (GAN) and Convolutional Neural Networks (CNN) can commonly be used in the field of image processing in fog.
V-A3 Edge
The concept of edge computing is relative to cloud computing. The processing method of cloud computing is to upload all data to the cloud data center or server in the computing resource set for processing. Any request to access this information must be submitted to the cloud for processing [63]. Edge computing is a computing model that moves resources and provides edge intelligent services at the network edge near objects or data sources, by which improve service quality and information security. In brief, edge computing analyzes the data produced by the terminal directly in the local device or network near the data generation, without transmitting the data to the cloud data processing center [47]. Compared with cloud computing, edge computing has shorter network latency, less resource consumption, and higher security of the data to the cloud data processing center.
The so-called edge AI combines two emerging technologies: edge computing and AI. However, the implementation of edge computing is based on the same basic premise, that is, generating, collecting, storing, processing and managing data locally rather than in remote data centers. Edge AI improves this concept to the device level, using ML to imitate human reasoning to reach user interaction points, such as computers, edge servers or IoT devices [48]. Typically, these devices can operate without an Internet connection and make decisions independently. Well-known examples of edge AI technologies include virtual assistants, such as GPT-4o, Apple’s Siri, or Amazon Alexa. When the user says ”hey”, these tools will recognize and learn what the user is saying (i.e. ML), interact with cloud-based application programming interfaces (APIs), and store the learned knowledge locally.
V-B Application Architecture
Monolithic and microservice, as the two main application architecture patterns [73], each has unique advantages and application scenarios that can support AI applications and models.
V-B1 Monolithic
Monolithic Architecture, was the traditional model for software development and deployment, used in the past by large companies. And its functionality is encapsulated into one single application. The advantage of this architectural pattern is that it is much easier to develop, deploy, debug and monitor [74]. Because the interaction between components is directly completed via memory, the performance is better. For some simple and initial AI systems, or applications with small business scales and infrequent changes in requirements, monolithic architecture may be a suitable choice.
However, with the arising of the business scale and demand, MA has encountered many challenges and shown obvious shortcomings. Due to its inherent tight integration, this design pattern has limitations in scalability, adaptability, and ease of maintenance. Whenever specific modules need to be extended, the entire application needs to be recompiled and deployed, which can lead to low resource efficiency and waste [20]. In addition, as the code repository grows, the complexity of development, testing, and deployment also increases, thereby increasing maintenance costs and the risk of errors. Therefore, in large and complex applications, developers are seeking more flexible and scalable architectural patterns, such as microservices architecture, as introduced in the next section.
V-B2 Microservice
Microservice, as an emerging distributed system architecture model, is gradually changing the landscape of software development. The core idea is, unlike monolithic, to decompose complex large-scale applications into small and independent service units, each focusing on specific business functions or integration domains. Considering the fact that microservices architecture is composed of multiple small components, iterative upgrades of applications are more flexible and efficient [20]. Especially for large and complex AI systems, such as e-commerce platforms, social media platforms, etc., microservice architecture is a more suitable choice. These applications typically consist of multiple subsystems, each of which can be independently developed. For example, Taobao has dozens of independent systems, all of which are typical microservice architectures that can support rapid business development and iteration.
Even so, with the widespread application of microservice architecture, it also faces challenges such as service governance, network transmission efficiency, service expansion, and version iteration. Nevertheless, microservice architecture remains an important development direction for enterprise IT architecture, and its potential and advantages cannot be ignored [75]. Future research will require an in-depth exploration of how to overcome these challenges and further promote the development and application of microservice architectures.
V-C IoT Use Cases
IoT is truly revolutionary, presenting an extensive array of diverse and impactful use cases that are reshaping numerous aspects of our modern world. We can classify IoT use cases into two main categories: static and mobile. As the name suggests, static use cases are like agricultural monitors, which are fixed and usually do not need to be moved, while mobile use cases such as in-car telemetries and wearable devices may frequently move [76].
V-C1 Static
Static IoT user cases based on edge AI technology are gradually becoming a key driving force for digital transformation. In these cases, data collected by static IoT devices (such as cameras, sensors, etc.) is directly analyzed in real-time on edge devices, and rapid decision-making and response are achieved through edge AI [75]. In the field of agricultural ecological environment monitoring, the agricultural IoT predominantly employs high-tech approaches to establish sophisticated agricultural ecological environment monitoring networks, and uses wireless sensor technology, information fusion transmission technology, and intelligent analysis technology to perceive changes in the ecological environment [4]. In 2002, researchers at the University of California, Berkeley conducted a 9-month periodic environmental monitoring of the habitat of the Shanghai Swallow on Duck Island using wireless sensor networks. Regional static MICA sensor nodes were deployed to achieve unmanned and non-destructive monitoring of sensitive wildlife and their habitats. Some countries, including the United States, France, and Japan, have primarily focused on integrating the establishment of agricultural information platforms covering the whole country to achieve automatic surveillance of the agricultural ecological environment and ensure its sustainable development of the agricultural ecological environment.
V-C2 Mobile
The emergence of emerging AI such as DL has brought new innovations to mobile animal networking. In the field of mobile IoT, edge AI achieves real-time and efficient data processing by deploying the computing power of AI at the edge of devices. For example, in smart health monitoring applications, edge AI enables wearable devices to analyze user physiological data in real-time, providing real-time health feedback to users without uploading data to the ”cloud” for processing [77]. These devices harness DL technologies to scrutinize user health metrics and furnish guidance for subsequent lifestyle modifications, yielding substantial benefits, particularly for infants, young children, and the elderly. Due to data not being uploaded to the cloud, it can effectively protect user privacy and security. In addition, in the field of intelligent transportation, AI technology provides timely guidance for vehicle operation by analyzing local data. Similarly, many frameworks utilize DL techniques to predict parking occupancy rates, reduce the time required for vehicles to search for parking spaces, and improve urban traffic management [6]. These user cases demonstrate the enormous potential of edge AI technology in improving the performance of mobile IoT applications and enhancing user experience.
V-D Methods
In this part, we discuss the dominant methods employed in edge AI which include heuristic algorithm, meta-heuristic algorithm, ML and Deep Reinforcement Learning (DRL). These four types of algorithms have their own characteristics and application scenarios, aiming to optimize the accuracy and performance of AI models and applications.
V-D1 Heuristic
People often refer to methods inspired by the laws of nature or the experiences and rules of specific problems as heuristic algorithms. The current heuristic algorithms are not entirely based on natural laws, but also come from human accumulated work experience. They supply a practical solution for each instance of the combinatorial problem to be solved and maintaining an acceptable cost in terms of computation time and space, and the degree of deviation between the practical solution and the optimal solution may not be ascertainable before the implementation in advance [60]. Heuristic algorithms are a technique that enables the search for the best possible solution within an acceptable computational cost, but may not necessarily guarantee that the obtained solution is the best scheme. In most cases, it is impossible to describe the degree of approximation between the obtained solution and the optimal solution.
In the field of edge AI, heuristic algorithms have been widely used because they conform to the characteristics of human cognitive thinking. Usually, edge devices have limited computing resources [78], so reasonable resource scheduling is particularly important. Heuristic algorithms can find satisfactory solutions in a relatively large search space within a short amount of time, so they can help our applications deploy to suitable nodes. In addition, heuristic algorithms can perform efficient data analysis on edge devices, greatly reducing dependence on cloud computing resources and lowering network response latency [79].
V-D2 Meta-Heuristic
Meta heuristic is a computational intelligence-based mechanism used to solve complex optimization problems for optimal or satisfactory solutions. The meta heuristic algorithm obtains a sufficiently good solution by searching the space [48]. Meta heuristic algorithms can be seen as an algorithmic framework that can be applied to different optimization problems with slight modifications. In edge AI, the application of meta heuristic mainly focuses on two aspects:
Model optimization: The reasonable deployment of AI applications in edge devices has always been a headache inducing issue. AI applications are computationally intensive services that require high computing resources, while computing resources in edge devices are usually limited [65]. The fully heuristic algorithm can help find the optimal parameter configuration for AI models, so that these AI applications can maintain high performance while occupying as few resources as possible.
Resource management: In edge AI systems, multiple tasks may need to run at the same time and they need to share limited resources [80]. The meta-heuristic algorithm can optimize the allocation of these resources, ensuring that each task receives sufficient resources to run efficiently without depleting the entire system’s resources.
V-D3 Machine Learning
As is well known, ML technique forms the foundation of AI and plays a pivotal role in many applications, such as recommendation systems, text generation, and so on. As edge AI gains prominence, ML is increasingly finding its way onto devices located on the fringes of the network. These edge devices prioritize data processing close to its source, aiming to minimize transmission delays, accelerate response times, and alleviate the reliance on centralized servers [6]. Consequently, edge AI systems are often tasked with handling vast volumes of data in real-time scenarios. Following certain ML algorithms, such as sophisticated DL models, these systems can swiftly analyze and process data at the edge, extract crucial insights, and empower a diverse array of real-time applications encompassing autonomous vehicle operations, intelligent manufacturing processes, and robust security systems. In addition, ML techniques can be employed to detect anomalies in real-time on edge devices, such as detecting product quality issues during the manufacturing process, or detecting changes in patient health status in the medical field [81].
V-D4 Deep Reinforcement Learning (DRL)
Many application problems in AI require algorithms to make decisions and execute actions at every moment. For Go, each step requires determining where to place the pieces on the chessboard in order to defeat the opponent as much as possible; For autonomous driving algorithms, it is necessary to determine the current driving strategy based on road conditions to ensure safe driving to the destination; This type of problem has a common characteristic: to make decisions and actions according to current conditions in order to achieve a certain expected goal [82]. The ML algorithm used to solve such problems is called reinforcement learning (RL). Although traditional reinforcement learning theories have been continuously improved in the past few decades, they are still difficult to solve complex problems in the real world.
DRL is a type of DL with reinforcement learning methods, which enables models to have stronger learning abilities, indicating that machines can autonomously understand and learn the human visual world. Simply put, just like humans, this means inputting visual and other perceptual information, and then directly outputting actions through deep neural networks without the need for manual production. As introduced before, DRL can solve specific problems in edge devices, such as in car systems where the device perceives the surrounding environment and road conditions on its own without the need for human intervention, and selects the appropriate driving route [83]. In addition, DRL has also driven the development of other fields, such as smart homes. [4]
V-E Resource Management
Edge computing, which deploys computing, storage, network, and other resources at the network’s edge, can drastically minimize data transmission delay, enhance data processing efficiency, and relieve bandwidth demand on the core network [4]. However, with the growing number of edge devices and the complexity of applications, how to efficiently manage these resources has become an urgent problem that must be addressed. This section will review how to make resource management in edge environment from the aspects of resource provisioning, resource allocation, application placement, and workload distribution and prediction.
V-E1 Provisioning
As the name suggests, resource provision is the way of provision of resources by resource suppliers based on the users’ pre-established needs and supply strategies. This type of strategy is usually divided into two types: dynamic and static strategies. Static strategies are determined based on user resource needs and constraints, such as QoS and SLAs [4, 84, 85]. Static strategies are more suitable for stable workloads. For applications with large fluctuations in resource demand, we usually use heuristic algorithms and ML algorithms to predict in advance, which are called dynamic strategies.
V-E2 Resource Allocation
Edge AI is an offline service during model training, different from traditional online services such as microservices, web applications, and API services. Sometimes, it is not sensitive to latency, but usually requires higher resources such as GPU and memory. The computing power of edge devices is often limited by their limited resources, which are often unable to support or complete computationally intensive tasks, such as model training, within an acceptable deadline. This approach at this point is to offload these offline tasks and transfer them to edge servers for completion [1]. Edge devices are used for services that need to meet low latency, such as web services and human-computer interaction. Of course, we can also combine some DL techniques to achieve more reasonable resource allocation. This approach aims to minimize resource waste and enhance application performance.
V-E3 Application Placement
Application placement is also an indispensable factor to consider in edge computing resource management, which means formulating an assignment of applications to servers that maximize the QoS for all users in order to optimize the performance for some components that are sensitive to latency such as interactive online games, face recognition, etc. Through reasonable application placement, such as some AI-based approaches [86], the computing and storage resources of edge nodes can be fully utilized, improving resource utilization and efficiency.
V-E4 Workload Distribution and Prediction
Generally, the infrastructure architecture where our application can be described as three separate layers: cloud layer, edge layer, and IoT layer. The workload of the application will have its own characteristics distributed in these three layers  [87]. The cloud layer is located at the top level of the architecture and The cloud layer, situated at the top of the architecture, is a robust cluster comprising thousands of virtual machines. Deploying large-scale AI models on the cloud layer is a good choice. The edge layer is located between the cloud layer and the IoT layer composed of a set of nodes that deploy several devices like routers or switches, and it is responsible for load-balancing traffic from the cloud layer and aggregating and analyzing data from the IoT layer. We can deploy some web servers and small-scale AI applications on this layer. The IoT layer is the source of data in this architecture, mainly composed of sensors and wireless devices, such as temperature sensors, cameras, and Bluetooth [78]. These devices need to have high sensitivity and feedback very quickly to user instructions.
Furthermore, accurate predictions can lead to more rational resource management. This is beneficial for both users and service providers. Previous research and technologies such as ARIMA [88] and Holt Winter[89] were based on linear temporal prediction. However, these technologies often exhibit poor prediction accuracy. With the rise of DL, technologies such as neural networks have been widely applied in data prediction, such as weather, transportation, and finance systems. Especially for recurrent neural networks (RNN), their inputs not only focus on current data, but also contain information from a period of time in the past. Therefore, it is often used to predict time-series related data. Using recurrent neural networks to predict workload in edge computing is a promising approach worthy of consideration.
V-F ML Model Sizing
In the edge computing scenario, the size of the model becomes the key factor to determine whether it can be successfully deployed and applied. Especially in intelligent camera monitoring systems, due to the limitations of edge devices in computing power, storage space, and energy supply, using a full model size DL model is often impractical [90]. To overcome these limitations, we usually adopt a strategy of reducing model size. In this section, we analyze and compare the two different strategies of model sizing.
V-F1 Reduced
The training cost and efficiency of AI models are important metrics to assess the quality of the model. Nowadays, many AI giants are progressively increasing the size of model parameters and the volume of training data. The model parameters of GPT-3.5 have reached 175 billion [91]. While this approach significantly enhances model accuracy, it also markedly escalates training costs and hardware requirements, necessitating a trade-off between accuracy and cost. In this way, we need to strike a balance between accuracy and cost. In recent years, there have been numerous studies in this area, such as DenseNet [92], EfficientNet [93] and EfficientNetV2 [94]. The goal of these works is to train models to achieve satisfactory accuracy with fewer model parameters. In the field of edge computing, where hardware resources are limited, the reduced size model will certainly become an important trend of edge AI in the future.
Model pruning and model quantization are well-established methodologies for achieving model size reduction. However, nowadays, how to design a lightweight and high-precision neural network has become a focal point of research in the field of AI, such as MobileNet [95] and ShuffleNet [96]. These models have small parameters and high computational complexity, making them very suitable for running on edge devices.
V-F2 Full
Unlike MobileNet and ShuffleNet, GPT-3 is a full model with 175 billion parameters, which is hundreds of times the number of GPT-2 parameters (3 billion). Tom Brown [97] demonstrated that GPT-3 has completed various NLP tasks, such as translation, question answering, etc., with minimal sample training. Due to its outstanding performance in the domain of NLP, this model has greatly promoted the development of large language models. Currently, many edge computing frameworks, such as KubeEdge [98], have integrated plugins that support the deployment of these extensive language models, thereby extending their applicability and utility in edge environments.
V-G Heterogeneity
Heterogenous environments in edge devices are employed to run various IoT applications. Their diversities are embodied in three aspects: computational heterogeneity, hardware heterogeneity and platform heterogeneity.
V-G1 Computational
Computational heterogeneity in edge computing emphasizes the variability in application behavior during computational operations. For AI applications, there is a large amount of vector operation logic in the model code, which determines that such applications are suitable for parallel computing rather than serial computing. For web services, universal computing is the main approach [99]. This distinction manifests in hardware requirements, where AI applications rely on GPU acceleration, whereas web services operate efficiently with CPU resources alone.
For many microservices, such as the web services mentioned earlier, their performance bottleneck often is not in CPU but in disk read and write speed, as most of the time is spent accessing databases. In other words, they are IO intensive services rather than computationally intensive. In order to reduce network latency, microservices are usually deployed in edge nodes. How to reduce the performance loss caused by slow disk read and write speeds is a problem we need to consider.
V-G2 Hardware
Edge devices have many differences in processor and hardware architecture due to the computing characteristics of the applications deployed on them. The instruction set of CPU can be divided into two categories: ARM and AMD, and software running on different instruction sets may have differences in performance. Some infrastructure deployed on edge nodes, such as routers and switches, are responsible for tasks such as data forwarding and protocol conversion, and therefore require CPU support. The application of image generation and virtual reality requires high-performance graphics rendering and environment recognition. In addition to processing CPUs, GPUs and other chips are also necessary. Nowadays, many container frameworks have support for CPU and GPU hardware resource isolation, such as Docker and Container, which can eliminate the impact of service instability caused by hardware resource competition.
In 2018, Google launched Edge TPUs [100], specially designed for inference and training of neural networks on edge devices with limited resources. Edge TPUs demonstrate strong capabilities in computer vision [101]. Some IoT applications for autonomous driving and facial recognition can benefit greatly
V-G3 Platform
Due to the rise of edge computing, the world’s major technology giants have also launched their own edge computing platforms. For example, Amazon’s AWS IoT Greengrass, Microsoft’s Azure IoT Edge, and Google’s Cloud IoT Edge. They all support the effective operation of AI models on edge devices, providing service management and data analysis capabilities. Other open source platforms also deserve attention, such as KubeEdge and OpenYurt, which are extensions of Kubernetes in the field of edge computing and provide container management, automatic operation and maintenance and other functions.
V-H Security
With the rapid development of edge computing techniques, an increasing number of enterprises and organizations are deploying edge computing solutions to meet high demands for real-time capabilities, security, and privacy protection. However, simultaneously, the edge computing environment also faces numerous security challenges [6, 102]. To ensure the stable operation of edge computing systems and data security, we need to consider and ensure security comprehensively from three aspects: Platform, Host, and Network.
V-H1 Platform
Blockchain is a distributed, decentralized, and tamper proof database. It is often used to build a secure and trusted intelligent platform, which can solve the security problems in edge computing. Zhang et al. [103] utilized blockchain technology to construct a highly secure trusted edge platform, providing a secure environment for AI applications on edge nodes. Wang et al.  [104] proposed an integrated trust evaluation mechanism based on cloud and edge computing, along with a new architecture of service templates and balanced dynamics, to address security challenges. In this architecture, the design of edge networks and edge platforms is aimed at reducing resource consumption and ensuring the scalability of trust evaluation mechanisms, respectively. Other security technologies such as Role Based Access Control (RBAC) have also been widely applied to some distributed platforms, such as Kubernetes.
V-H2 Host
Host security is defined as the security of all hardware and software deployed on a single edge server or device. Due to the proximity of edge devices to the human body, such as healthcare systems and intelligent driving systems. Imagine that if a car is using intelligent driving and its intelligent driving system is hacked, it will pose a serious threat to the safety of passengers and other vehicles on the road [105].
We can take many measures to defend against external attacks on the host. Firewall rules can be configured to block access from unauthorized IP addresses. Moreover, by installing antivirus software and regularly updating patches, the security factor can also be improved.
V-H3 Network
Distributed Denial of Service (DDoS) attack, which causes significant economic losses to society every year, is one of the most common attack methods in computer networks, and it also has strong destructive power on IoT devices. From the time of the 2016 botnet Mirai attack on KrebsOnSecurity [106] and Dyn [107], it can be seen that DDoS attacks are seriously threatening the security of IoT applications. With the development of edge computing, the threat of such attacks to large-scale IoT devices is growing, which may lead to incalculable economic losses. For example, in the field of automation, AI technology is widely used to make decisions and adjust plans. If edge AI is subjected to network attacks, it can lead to AI models making incorrect decisions, resulting in product quality issues.
Although edge nodes exhibit the potential to isolate most of the IoT data at the network edge and detect and intercept attacks near the source in the first place, they encounter significant challenges in practical applications. The main reason is that edge nodes are unable to capture the aggregated network traffic required for IoT DDoS detection, nor can they scale and provide the necessary resources like elastic clouds [108]. Therefore, directly deploying existing cloud based defense solutions on edge nodes is far from achieving ideal results. We need to redesign the DDoS defense scheme based on edge computing to solve the special and severe security problems in the edge environment.
V-I Scheduling
Resource scheduling is the process of efficiently allocating and managing system resources, ensuring optimal utilization of resources according to demand and priority. In edge environments, resource scheduling is exceedingly crucial for achieving real-time, low-latency services. Especially in AI scenarios that have high demands for computing and network resources, only by allocating edge device resources reasonably can we support the rapid response and efficient operation of AI applications, and improve overall system performance and user experience. Resource scheduling is also a hot research direction, and there has been a lot of work in this area before [4, 102, 6]. In this section, we discuss scheduling from the following four granularities, because the four constitute the core unit for application deployment and management in container orchestration systems such as Kubernetes and KubeEdge.
V-I1 Container
Containers are a software virtualization technology that laid the foundation for the development of microservices. Nowadays, containers are also widely used in the field of edge computing.There are also many researches on containers in edge computing, which stems from the growing demand of users for millisecond delay computing. In  [109], the authors elucidate the concepts of container placement and migration between edge servers, and proposes a container scheduling framework grounded in multi-objective optimization models or graph network models.
In addition, some open-source container orchestration and scheduling frameworks are worth paying attention to, such as KubeEdge. KubeEdge can extend its powerful cloud computing capabilities to edge devices. Especially suitable for some AI applications, model training can be completed in the cloud and then deployed to edge devices. In addition, KubeEdge can optimize scheduling performance based on different AI application business scenarios by configuring the algorithm and parameters of kube-scheduler.
V-I2 Task
In the scheduling task of edge computing, there are usually two problems to be solved: scheduling time and resource allocation. Most previous research [110, 111, 112] on task scheduling has focused on these two aspects. With the advancement of AI technology, ML technology has shown unique advantages in task scheduling. Markov Decision Process (MDP) is an robust and effective method for modeling temporal data and providing high-precision predictions. The problem of resource allocation in edge devices can be described as MDP, and the deep Q network (DQN) algorithm uses multiple replay memories to minimize the total delay and resource utilization. the study in [113] addresses the intricate issue of joint task offloading and resource allocation problem for computationally intensive tasks in fog computing. This intricate problem is formulated as a partially observable MDP, and the Deep Recursive Q-Network (DRQN) algorithm is adopted to approximate the optimal value function.
V-I3 Pod
In container orchestration systems such as Kubernetes, Pod is the smallest unit of work composed of several containers. Pod scheduling is the process of assigning Pods to a node based on a certain algorithm strategy, which is of great significance for ensuring high availability, resource utilization, and performance of the system. Pod scheduling is mainly controlled by kube-scheduler, and its process includes two stages: screening and scoring [114]. During the filtering phase, the scheduler checks all nodes to determine which ones have the resources (such as CPU and memory) and other requirements (such as node selector labels) needed to run Pod. Then, the selected node will enter the scoring stage, and the scheduler will rate each node based on a series of criteria such as node affinity, resource utilization, etc. The node scoring highest will be designated as the running location for Pod. kube-scheduler supports custom scheduling plugins, and users can develop some extension plugins based on the business characteristics of the enterprise.
V-I4 Service
Service refers to software or system components deployed at the edge of a network that provides specific functions or resources to meet the real-time, low latency, and high bandwidth needs of users or devices. Service can be a computing service, data processing service, storage service, or any form of network service that optimizes resource utilization and reduces data transmission latency, bringing better service quality and experience to users.
Service scheduling is the deployment, allocation, and scheduling process for these services [38]. In the edge computing environment, Service Scheduling is responsible for arranging and scheduling the execution sequence and location of services reasonably based on application requirements, resource conditions, and network conditions. Effective service scheduling can ensure that the service can efficiently use limited edge computing resources, achieve load balancing, reduce service latency, and improve the performance and reliability of the entire system.
V-J Container Migration
In distributed and cloud computing environments, containers need to be migrated from one node to another due to node failures, load imbalance, resource upgrades, and other reasons. At this point, container migration technology is needed to achieve rapid migration and recovery of containers.
V-J1 Stateful vs Stateless containers
Stateful containers and stateless containers are two major classifications of containers, which are important criteria for container expansion, contraction, and migration.
Stateful container: The so-called state essentially refers to the data in the running container. When migrating such containers, it is usually necessary to migrate their data together, such as a database. Due to the involvement of data replication, such containers need to consider issues such as data loss and data integrity. Specific migration tools or strategies may be needed to ensure accurate migration and recovery of data [115]. All containers managed by a StatefulSet controller in Kubernetes are considered as stateful.
Stateless containers: These containers are containers that do not save any state during runtime. For example, a web server that provides services for static pages, treats each request as independent, and the container does not need to remember previous interactions. The migration process is very simple, just pull up the container on other nodes and delete the container from the original node. All Pods under a Deployment in Kubernetes are stateless.
V-J2 Inter versus Intra cluster migrations
Inter-cluster migration: When a company or organization needs to migrate its data center from one geographic location to another, inter-cluster migration is an indispensable step. Inter-cluster migrations involve node migration between different clusters, typically requiring consideration of cross-cluster communication factors such as network latency and bandwidth limitations [116]. Due to the collaborative work of multiple clusters and nodes involved in cross-cluster migration, the migration process is relatively complex and requires ensuring data consistency and service continuity.
Intra-cluster migration: In a cluster, migration within the cluster can take effect when a node experiences performance degradation or longer response time due to excessive workload. Administrators or automation tools can migrate a portion of the workload (such as containers, virtual machines, or services) on that node to other nodes in the cluster to balance the load and optimize performance [42]. Compared to inter-cluster migrations, The complexity of intra-cluster migrations is relatively low because it only involves nodes and data migration within the same cluster.
V-J3 Migrations at cloud/edge/fog
Migration at cloud is the process of migrating applications, data, and other business processes from traditional local devices or servers to cloud platforms, including the migration to IaaS, PaaS, and SaaS [117]. IaaS migration is the most ideal and applicable cloud migration solution. Because we can entrust all programs and data to cloud vendors such as Alibaba Cloud and AWS. Users do not need to consider all operational and deployment issues.
Edge computing has become an important technical support for the development of IoT [118]. A thorny problem in edge computing is service migration, especially in the mobile IoT device environment. Due to the limited coverage of a single edge server network, the migration of mobile services between servers is likely to reduce the QoS of the services. State preservation of services (such as stateful services), data loss, and cost control have become challenges in the migration of edge computing services.
Migration at fog is the process of migrating applications, services, or data from traditional centralized data centers or cloud environments to a fog computing environment. The purpose of this migration is to achieve low latency, bandwidth optimization, enhanced security, improved scalability, and fault tolerance. Fog migration involves redesigning applications to adapt to the distributed architecture of fog computing, including modular design and the ability to handle network dynamics [45, 119]. Fog migration can provide more effective support for IoT devices, mobile devices and other applications that need rapid response, and achieve the goal of intelligent edge computing.
V-J4 Simulations versus real-world testbed migrations
When discussing container migration, two different testing and validation methods are usually involved: simulations and real-world testbed migrations. Here is a comparison between these two methods:
Simulations: It uses models to replace actual or conceptual systems for training, analysis, argumentation, experimentation, experimentation, and planning methods, techniques, and activities [4]. Simulations can predict system performance and efficiency, validate and iterate modeling and simulation through real experimental data, support, optimize and expand experimental identification, accelerate development and reduce risk costs [60].
Real world testbed migrations: Its definition is the process of testing and validating a system or application in a real physical environment, involving the migration of the system or application from one environment to another. Since it is conducted in a real-world environment, it can directly evaluate the performance, reliability, and safety of the system or application under actual operating conditions to ensure that it meets practical needs [1].
When conducting container migration testing and validation, simulations and real-world testing platforms are usually combined. Simulation can quickly validate concepts and strategies in the early stages, while real-world testing platforms are used to test and optimize migration strategies under conditions close to actual operational environments. This combination of methods can balance cost, time, and accuracy of results, providing a comprehensive evaluation for fog migration.
V-K Container Scaling
With the continuous development of cloud computing and container technology, container scaling has become an important means to ensure application performance, high availability, and resource optimization. This section will explore the strategies and practices of container scaling from two key perspectives: firstly, the scaling decisions of proactive and reactive, which exhibit different characteristics and advantages in dealing with load changes; next horizontal vertical and hybrid scaling strategies represent how effectively adjust container resources in different scenarios.
V-K1 Proactive versus reactive scaling decisions
The scaling decisions of Proactive and Reactive reflect two different strategies, which have a significant impact on the performance and resource allocation of container applications. The following are specific explanations of these two strategies:
Proactive scaling decision: This method will use historical data of container load to train a specific AI model, through which future changes in container resource load can be perceived and predicted in advance [120]. It allows administrators or systems to automatically adjust resources to maintain optimal performance and efficiency. For example, this strategy can predict based on historical data that as long as it reaches 7pm or 8pm, the QPS of AI applications will significantly increase because everyone is off work, which is the entertainment time at night.
Reactive scaling decision: This is a strategy that utilizes third-party resource monitoring tools, such as Promethues [121], to make real-time decisions on the number of replicas and resource allocation in containers. The container orchestration tool determines whether to expand or dissolve based on the resource change data of the relevant containers in the monitoring tool. When the load increases, reactive scaling will start adding containers; When the load decreases, it will decrease the number of containers [122]. The decision-making of reactive scaling is based on real-time load data. When training the model, the utilization of GPU and GPU memory inside the container may reach 80%-90%. At this time, the system will immediately detect the high utilization rate and scale up the capacity promptly.
V-K2 Horizontal, Vertical and Hybrid scaling
Three types of scaling techniques are described below:
Horizontal scaling: It is a way to cope with load changes by increasing or decreasing the number of container instances [123, 124] (such as Pods, container groups, etc.). It can respond very quickly to load changes and adjust overall processing power by adding or removing container instances. Each container instance is independent and has good fault isolation, a fault in one instance will not affect other instances. Horizontal scaling is very suitable for scenarios with stateless services and the need to handle a large number of concurrent requests.
When AI applications need to handle a large number of concurrent requests and each request has a relatively short processing time, horizontal scaling is a good choice. For example, online recommendation systems, real-time advertising delivery systems, etc. In some application scenarios that require a large amount of computing resources (such as CPU, GPU, memory, etc.), horizontal scaling can provide sufficient resources by adding more machines. For example, DL model training, large-scale image recognition, etc.
Vertical scaling: It is adjusting the processing power of a single container instance by increasing or decreasing its resource allocation (such as CPU, memory, storage, etc.). It does not require managing multiple container instances, only adjusting the resource allocation of a single instance and can accurately adjust resource allocation based on actual load conditions, avoiding resource waste. Vertical scaling is suitable for stateful services[125]. However, this scaling strategy also has its drawbacks, as it poses a challenge to the computing and storage capabilities of individual machines.
When AI applications encounter performance bottlenecks stemming from the capabilities of individual nodes, vertical scaling offers an effective solution by enhancing hardware capabilities, such as deploying faster CPUs, increasing memory capacity, or leveraging more efficient GPUs. For AI applications that do not necessitate extensive concurrent processing or substantial computing resources, vertical scaling can serve as a more cost-effective and straightforward approach, minimizing complexity while maximizing performance within the confines of a single node.
Hybrid scaling: Hybrid scaling is a scaling strategy that combines both horizontal scaling and vertical scaling [84]. Based on the load characteristics and requirements of the application, use both horizontal and vertical scaling methods to optimize resource allocation and performance. Being able to flexibly choose scaling methods based on different scenarios and needs, and combining horizontal and vertical methods can more effectively utilize resources, and improve application performance and stability.
For AI applications where demand often changes or is difficult to predict, hybrid expansion can dynamically adjust the ratio of horizontal and vertical expansion based on actual demand.
VI Comparisons of Existing Edge AI approaches based on Taxonomy
In this section, we compare the existing edge AI approaches based on the proposed taxonomy.
VI-A Infrastructure
Cloud computing, fog computing, and edge computing play different roles in realizing offline, low-latency, privacy-preserving AI services [126]. Among them, cloud computing provides powerful computing and storage resources for training large-scale DL or other algorithm models, and processing massive amounts of data, which are usually used in Edge AI to handle time-insensitive tasks, such as large model training, multi-data analysis and model optimization, and finally, cloud computing distributes well-trained models to various user devices;
Fog computing moves computing resources to the edge of the network, reduces data transmission latency, improves response speed, and is typically used in Edge AI to handle tasks that require high real-time responses, such as speech recognition.
Edge computing deploys computing resources directly near terminal devices to further reduce data transmission latency, which is usually used for real-time reasoning and decision-making in Edge AI, such as intelligent monitoring, smart home, etc., edge computing can realize real-time processing of data on user devices, maximize the protection of users’ data privacy, and at the same time reduce the dependence on network bandwidth and reduce the pressure on the core network.
In summary, cloud computing, fog computing, and edge computing have their own focus in Edge AI, and these three together build a complete edge intelligence ecosystem. Cloud computing provides powerful computing and storage resources, fog computing emphasizes real-time and low latency at the edge of the network, and edge computing enables real-time data processing and decision-making closest to the end device. These three work together to provide comprehensive support for the development of AI at the edge.
As shown in Fig. 8, we compare the emphasis, advantages and disadvantages of cloud computing, fog computing, and edge computing under different indicators:
Figure 8: Advantages, disadvantages and emphases of the three computational paradigms
VI-B Application
Monolithic and microservices have their own advantages and disadvantages in edge AI, and we will compare the advantages and disadvantages of these two in terms of flexibility, performance and resource utilization, as well as deployment and scalability.
flexibility:Since all functional modules of the monolithic architecture run inside the same application, if we need to modify a module, we may have to recompile and deploy the entire application, so the monolithic architecture is not conducive to modularity and independent development, that is, it is less flexible; Each microservice in the corresponding microservices architecture can be deployed, scaled, and updated independently, making it easy to develop and maintain independently, thus increasing the flexibility of the overall application.
performance and resource utilization:Monolithic architectures may have resource contention and performance bottlenecks because all modules share the same process and resources, but from a resource utilization perspective, monolithic architectures may make more efficient use of resources because they do not require additional communication and management overhead; The microservices architecture, on the other hand, can independently deploy and scale the corresponding microservices according to the needs, thereby improving the performance of the application. From the perspective of resource utilization, services in a microservice fabric need to communicate with each other, which may increase the latency and bandwidth consumption of the system, and the microservice system may require more resources to manage and run due to the need to maintain multiple services.
deployment and scalability:Monolithic architectures are typically simple and easy to implement and deploy, but they often lack scalability to cope with frequently changing requirements; The corresponding microservices architecture, while more complex to deploy and often requires additional development and management efforts, scales flexibly and allows services to be added or removed quickly as needed.
In summary, the monolithic architecture focuses on simple deployment and performance optimization, which is suitable for simple, relatively fixed edge AI scenarios, while the microservice architecture focuses on flexible scaling and maintainability, and is suitable for complex edge AI application scenarios that need to be dynamically adjusted.
VI-C IoT Use Cases
There are significant differences between static and dynamic user use cases for IoT in edge AI. Static IoT user use cases typically involve the collection and analysis of static data, with relatively low data flow and frequency, low real-time requirements, and less variation in the location of the device. In contrast, dynamic IoT user use cases involve the collection and processing of real-time data, with higher data flow and frequency, higher real-time requirements, and the location of devices that can change frequently. In addition, static IoT use cases typically have relatively low energy consumption, while dynamic IoT use cases may consume more energy due to continuous data transmission and processing. Static IoT use cases tend to be more centralized when it comes to data processing and decision-making, while dynamic IoT use cases may require more distributed processing and decision-making. We give the specific differences between the two in Table IV:
TABLE IV: Comparison of IoT user use cases
Different use cases Static IoT user use cases Dynamic IoT user use cases
Data traffic Low High
Real-time Low High
Changes in location Low High
Energy consumption Low High
Data processing and decision-making Centralization Distributed
Figure 9: Comparison of different models
VI-D Methods
For the four main AI methods, heuristics[127], meta-heuristics[74], machine learning[128], and deep reinforcement learning[129], we will compare them from the perspectives of applicable scenarios and problem complexity, data scale and training cost, real-time requirements and resource consumption, and generalization. Fig. 9 shows the specific comparison:
In summary, choosing the right algorithm depends on the specific edge AI application scenario, data scale, data type, real-time requirements, and resource consumption. Heuristics and meta-heuristics are generally suitable for simple to medium-complexity problems, and the requirements for data and resources are generally not very high. ML and DRL are more suitable for dealing with some complex and nonlinear problems, and have high requirements on data volume, data quality, and computing resources.
VI-E Resource Management
With respect to the methods of resource provisioning, resource allocation, application placement, and workload distribution and prediction in edge AI resource management[130], we will further describe the relationship between these methods in detail, as shown in Fig. 10:
Figure 10: Process of Resource Management
As shown in Fig. 10, when deploying edge AI, we first need to ensure that the edge devices have sufficient computing, storage, and network resources, and once they have sufficient resource provision, we also need to allocate these resources to different applications or computing tasks. Resource allocation ensures that each application or compute task gets the resources it needs to meet its performance requirements. Once the resources are allocated, the application needs to be placed on the appropriate edge device. Application placement needs to take into account the characteristics and needs of each application, as well as the state information of the edge device to achieve the best placement strategy. Once the application is placed, we can send compute tasks and compute workloads to various edge devices, and the workload distribution enables parallel processing tasks and load balancing, thereby improving the performance and efficiency of the system. Finally, we can predict future resource demand and workload changes through ML models, for example, so that we can make adjustments and optimizations in advance.
VI-F ML Model Sizing
Regarding the deployment of AI models on edge devices, we generally have two deployment methods: compressed model [131, 132] and full model[133], and we compare these two methods from five aspects: model size, inference speed, accuracy, training and deployment cost, and application scenario, as shown in Table V:
TABLE V: Comparison of different deployment methods
Model Full model Compressed model
Model size Large Small
Inference speed Slow Fast
Accuracy High Slightly lower
Training and deployment cost High Low
Application scenario Resource-rich equipment Resource-constrained devices
VI-G Heterogeneity
The different types of heterogeneity[134] involved in edge AI deployment mainly include computing heterogeneity, hardware heterogeneity, and platform heterogeneity, as shown in Table VI:
TABLE VI: Comparison of different types of heterogeneity
Heterogeneity Computing heterogeneity Hardware heterogeneity Platform heterogeneity
Definition Different types of computing tasks Different kinds of hardware Devices with different functions
Example Image recognition; NLP CPU, GPU, FPGA, ASIC Cloud server; Edge device
Difference Differences in demand Diversification of hardware Differences between devices
VI-H Security
Starting from the security of deploying edge devices, we mainly consider platform security, host security, and network security[135]. As shown in Fig. 11, we will compare the roles of the three in detail:
Figure 11: Comparison of different types of security
VI-I Scheduling
For the resource scheduling categories in edge AI, there are mainly container scheduling, task scheduling, pod scheduling, and service scheduling [111], and we will compare these four different scheduling types from the aspects of emphasis, scheduling measures, and scheduling tools, as shown in Fig. 12.
Figure 12: Different types of resource scheduling methods
VI-J Container Migration
There are four main types of container migration[136, 137] in edge AI: stateful migration and stateless migration, intra-cluster migration and inter-cluster migration, cloud/fog/edge migration, virtual migration and real-world testbed migration. We have made a detailed comparison of the different migration methods, as shown in Fig. 13:
Figure 13: Different types of migrations
VI-K Container Scaling
Regarding container scaling[125] in edge AI, there are two main ways to actively scale and passively scale from the perspective of system response. From the expansion mode, there are mainly horizontal expansion, vertical expansion and hybrid expansion. As shown in Fig. 14, we make a detailed comparison of these two categories:
Figure 14: Different types of container scaling
Figure 15: Year wise Publication
VII Analysis and Result Outcomes
The survey enriches with various prospects associated with Edge empowered AI such as infrastructure support, IoT use cases, resource management strategies, security concerns and many more in the form of various state-of-the-art studies. The authors have systematically reviewed numerous articles in order to understand the prevailing the status of Edge AI in distinctive domains along with intelligent paradigms like ML and DL. There are lots of works going this direction to improve the life style of peoples and solve the real time problems. Hence, this section signifies the importance of our work referred in the form of year-wise papers, publication count, type of implementation (Simulation or experimental-based) and nevertheless QoS parameters addressed. Fig. 15 presents the year-wise analysis of related work carried in the form of number of papers referred from each year. The taxonomy of our study has been proposed with reference to articles from year 2015 to 2024. As depicted from Fig. 15, it is concluded that major chunk of the referred articles is recent and are from the year 2023. This clearly illustrates the fact that our survey includes the latest work done by the researchers.
Figure 16: Publication Statistics
Figure 17: Categorization of Articles
Figure 18: Summary of Future Research Directions
Apart from that, we rigorously reviewed the publication based statistics for the extensive study conducted highlighting its importance in real-time data processing. In total 1253 articles were collected during the data collection phase from various sources such as IEEE, ACM, Wiley, Science Direct, Taylor & Francis and Springer. Afterwards, filtering stage excluded collected articles based upon redundancy and inclusion and Exclusion criteria. The final stage comprises articles which the authors believe contributed the most towards shaping up the survey as depicted in Fig. 16 . Furthermore, the articles have been thoroughly reviewed and divided into 4 categories: review, Systematic Literature Review (SLR), implementation (simulation-based), implementation (Real/Testbed-based) and book. Fig. 17 illustrates that the major portion of the articles referred, are based on implementation (simulation-based), which signifies the fact that the implication of edge empowered-AI is yet to be tested on real-life IoT based use cases. Several real time work have been proposed by the researchers in recent years to improve the IoT applications based architecture using intelligent paradigm like ML, DL and reinforcement learning. This highlights a potential research direction for future studies to explore and validate edge-empowered AI in practical, real-world IoT environments. In addition, this article will motivate the researches to propose the novel solution to improve society 5.0.
VIII Future Research Directions
Edge AI is continuously evolving and showing potential across various domains, offering numerous opportunities for innovation and improvement. This section examines critical future research directions that promise to enhance the capabilities and applications of Edge AI. These directions include optimizing energy use, strengthening security, and integrating with next-generation networks like 6G, highlighting the transformative impact of Edge AI across multiple sectors as shown in Fig. 18.
VIII-A AI-driven Edge Data
AI-equipped edge devices can process data locally, providing real-time insights and actions without cloud transmission delays, crucial for applications like predictive maintenance in manufacturing. Using video, audio, and sensor data for anomaly detection, these devices enhance security in smart cities by identifying threats in real-time [138]. Additionally, AI-driven edge devices optimize energy usage by adjusting to weather, occupancy, and usage patterns in real-time [139]. Low latency and advanced privacy in AI model inference on the network edge benefit applications such as smart vehicles and drones [140]. The communication paradigm in Edge AI systems, designed to aggregate locally computed partial gradients at a central node, influences the performance of these edge nodes [140].
VIII-B Energy
Optimizing energy use in AIoT systems through intelligent edge computing need to focus on sophisticated algorithms for dynamic energy distribution and task scheduling, ML for workload management, and designing low-power hardware [141]. Improvements in communication protocols and integrating renewable energy sources will enhance efficiency and scalability. Literature [142, 143] reported that AI techniques for managing renewable energy sources has been investigated, emphasizing advanced ML models for accurate forecasting and optimizing energy storage and grid integration. Implementing edge AI will reduce latency and enable real-time decision-making, increasing system responsiveness and resilience [144]. It has been identified that incorporating edge AI in the Internet of Energy (IoE), highlighting secure edge computing, blockchain for data security, lightweight AI algorithms, standardization for interoperability, and 5G networks for low-latency communication [75, 145, 146]. Advances in federated edge AI and DRL will optimize energy distribution, enhancing the efficiency and resilience of IoE systems [147].
VIII-C Manufacturing
AI meets rising customer expectations for customization and high-value production by integrating capabilities at the network edges. It has been identified that integrating AI in manufacturing, enhancing collaboration with experts through tools like Google VisionAI for data science applications [148]. Key research areas include manufacturing scheduling and planning due to abundant data and productivity improvement opportunities. There is a need to investigate edge AI’s real-time analysis for predictive maintenance, quality control, and fault diagnosis in manufacturing, improving efficiency, reducing waste, and optimizing resources [149]. Implementing ML models at the edge allows continuous monitoring, early fault detection, and immediate corrective actions, enhancing intelligent manufacturing. Additionally, edge devices can optimize power generation and consumption by analyzing real-time data, promoting renewable energy use and cost savings.
VIII-D Smart Cities
The future of Edge AI in smart cities holds the promise of several significant advancements [150]. First, enhancing data processing capabilities at the edge reduces cloud load and latency, enabling real-time decision-making for applications like air and water quality monitoring, traffic management, and waste management. Integration with AI algorithms will facilitate more intelligent decision-making by analyzing sensor data from various city domains. Innovations in low-power hardware and efficient communication protocols are crucial to ensure scalability and energy efficiency. Developing secure edge computing and blockchain technologies will also address data security and privacy concerns [105]. Finally, advancements in 5G networks will provide the necessary infrastructure for high-speed, low-latency communication, further enhancing the responsiveness and resilience of intelligent city systems.
VIII-E Smart Transport
Applying DRL, specifically Deep Q-Networks (DQN), to mobile edge computing in smart transportation helps balance computing capability and traffic state. This approach highlights the need for further research on trade-offs and optimization techniques to enhance efficiency and performance in edge AI applications [151].
VIII-F Serverless Edge AI
Leveraging the flexibility and scalability of serverless architectures will significantly enhance the deployment of ML models in healthcare [116, 152]. This approach will enable real-time, cost-effective diagnostics without managing backend infrastructure. Future research will optimise neural network models for serverless environments, reduce cold start latencies, and enhance model performance through adaptive learning techniques [153]. Integrating serverless edge AI with IoT frameworks will facilitate continuous monitoring and rapid response in medical applications [154]. Additionally, improving interoperability and security protocols will ensure the safe and efficient handling of sensitive healthcare data in serverless architectures, paving the way for the next generation of intelligent, responsive, and secure healthcare systems.
VIII-G Quantum ML
Integrating Quantum Machine Learning (QML) with edge AI technologies, such as large, intelligent surfaces and visible light communications, will significantly reduce latency and enhance performance [155]. This combination enables efficient processing and decision-making at the network edge, which is crucial for managing the vast data generated by IoT devices and other edge sources [156]. Leveraging quantum computing capabilities in edge AI applications will achieve unprecedented network performance, leading to more responsive and adaptive AI systems [17]. The convergence of QML and edge AI in 6G networks will drive innovative solutions for real-time analytics and intelligent automation, meeting the increasing demands for low-latency and high-efficiency edge computing environments [157].
VIII-H Hardware
The physical boundaries for AI systems are set by the hardware of edge nodes which increases efforts in edge AI hardware design that is Nvidia’s Jetson TX2 is aimed at power-efficient embedded AI computing whereas Google’s Edge TPU is built for high-speed inference at the edge [20]. However, these devices mainly concentrate on handling entire tasks, especially local edge inference. Moving forward, we’ll see a variety of edge AI hardware designed specifically for different AI system architectures and applications [140].
VIII-I Heterogeneity
A multi-prototype-based federated learning enhances model inference by leveraging multiple weighted prototypes rather than relying on a single prototype, which can be incomplete and ambiguous [90]. This can aid in calculating these prototypes locally using clustering algorithms like k-means, ensuring effective representation of different client data distributions. Aggregating these multiple prototypes improves robustness against non-IID data distributions, leading to higher test accuracy and more efficient communication. This approach demonstrates significant improvements in accuracy and convergence rates, making it a promising direction for handling heterogeneity in Edge AI.
VIII-J Security
Utilizing AI-based quantum-safe cybersecurity automation, improving device and sensor security, and integrating quantum-safe solutions to counteract the threats posed by advanced quantum computing [158]. The research will focus on developing scalable and efficient cybersecurity systems, including AI-driven automation for threat detection and mitigation and using blockchain for secure communications [159]. Additionally, creating robust test environments for cybersecurity validation will ensure the effectiveness of these solutions in diverse operational scenarios [160]. These developments aim to provide a comprehensive security framework for future edge AI systems, ensuring resilience against evolving cyber threats.
VIII-K Privacy
Privacy enhancement in early health prediction through federated learning would be another interesting area to investigate [161]. Future directions in this field involve developing more advanced privacy-preserving techniques within federated learning frameworks to keep patient data secure during model training. Improvements in differential privacy and homomorphic encryption are essential for protecting sensitive health information [76]. Additionally, optimizing the communication efficiency between edge devices and central servers will mitigate privacy risks associated with data transmission. Integrating privacy-conscious AI models with real-time health monitoring systems, like wearable devices, can deliver immediate and secure health insights. Collaboration among healthcare providers, AI researchers, and policymakers is vital to creating standardized privacy protocols. Future research should also focus on scalable and adaptive federated learning methods capable of handling diverse and large-scale health data while maintaining high privacy standards.
VIII-L 6G and beyond
In the context of 6G and beyond, Edge AI is set to make several significant advancements; utilizing the ultra-low latency and high bandwidth of 6G networks will improve the deployment of AI models at the edge, facilitating real-time applications such as autonomous vehicles and smart cities [5]. The research will prioritize optimizing AI algorithms to meet the requirements of 6G, including dynamic resource allocation and energy efficiency. Combining quantum ML with 6G will enable more complex computations at the edge, enhancing predictive accuracy and decision-making processes. Additionally, improvements in secure edge computing and blockchain technology will address data privacy and security issues, ensuring robust and reliable edge AI systems. These advancements will collectively enhance the performance, scalability, and security of edge AI applications in a 6G environment [21].
IX Summary and Conclusions
The systematic review analysis of Edge AI provides a comprehensive overview of the present status of research in edge intelligence and its applications. The significance of these findings lies in the need for Edge AI systems to consider infrastructure, resource management, and the scale of ML models. According to the findings of the study, it is essential to conduct a thorough examination of both the positive and negative aspects of prior research in order to identify any potential gaps in research and to estimate prospective developments and concerns.
The study emphasises the importance of using a systematic approach to record and evaluate the existing research in the field of Edge AI. Moreover, it highlights the importance of implementing a standardised procedure to reduce the possible impact of discrepancies in the study. The results of this review have the capacity to ignite a new field of investigation in Edge AI and offer direction for prospective research in this field. The exhaustive examination of Edge AI offers profound insight into the most current study on edge intelligence and its realistic implementations. Moreover, this emphasises the importance of gaining additional understanding about the key factors that govern the choice of Edge AI infrastructure, as well as the effect of the scale of the model used for ML on efficiency and resource allocation.
To summarize, the comprehensive study on Edge AI is to fully evaluate the various AI methodologies. This study integrates all the feasible methodologies incorporates in edge intelligence or AI at the edge. This review is to examine the crucial factors that impact the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, and assess their impact on application efficacy and resource utilization. Furthermore, it investigates the influence of the size of a ML model on the efficacy and resource usage of an Edge AI application. A total of 78 studies have been chosen for this evaluation due to their specific focus on the application of AI in edge computing. In order to enhance our comprehension in the context of AI applied to edge computing, these research were categorized into multiple domains, including infrastructure, resource management, and ML model sizing, among others. Resource supply, allocation, scheduling, and job deployment are crucial considerations in the field of Edge AI.
Acknowledgements
M. Golec is supported by the Ministry of Education of the Turkish Republic. B. Ali is supported by the Ph.D. Scholarship at the Queen Mary University of London. H. Wu is supported by the National Natural Science Foundation of China (No. 62071327) and Tianjin Science and Technology Planning Project (No. 22ZYYYJC00020). F. Cuadrado has been supported by the HE ACES project (Grant No. 101093126). M. Xu is supported by the National Natural Science Foundation of China (No. 62102408), Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010251), Shenzhen Industrial Application Projects of undertaking the National key R & D Program of China (No. CJGJZD20210408091600002).